# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

### v1

@@ 7 Discussion and Conclusion
This work presented an extensive analysis of the robustness of segmentation models under localized corruptions, considering both natural and adversarial attacks. We first introduced metrics that extend the classic pixel-wise accuracy—commonly used in segmentation—to provide a more comprehensive analysis of spatial robustness, both within and beyond the corrupted regions. We then proposed a framework that enables a systematic and efficient evaluation of localized natural and adversarial corruptions by allowing the selection of key parameters, including corruption ratios and model patch sizes.
Our analysis of perturbations also improves upon existing strategies in the literature by demonstrating that, in dense prediction tasks such as semantic segmentation, targeting multiple pixels introduces a complex multi-objective problem. This complexity makes it difficult to assess the worst-case adversarial scenario using a single localized perturbation. To address this, we proposed region-aware multi-attack adversarial analysis, a method that applies multiple attacks while continuously adjusting the fooling area, thus providing a more realistic assessment of model robustness across different parts of the output.
All findings were evaluated across a diverse set of segmentation models in a driving scenario, specifically using the Cityscapes dataset [8]. Our results revealed distinct differences in model behavior under natural and adversarial corruptions. Furthermore, we highlighted the importance of finding a trade-off between these two robustness aspects, which are not necessarily correlated. As a preliminary step toward addressing this challenge, we explored ensemble strategies that combine convolutional-based models with transformer-based architectures.
For future work, we aim to deepen the analysis of such a trade-off also considering other tasks and theoretical insights. Another important direction is integrating the proposed localized robustness analysis into model training strategies, particularly through the lens of localized augmentation techniques (e.g., Cutout[11]) and CutMix[38], which are not commonly applied in complex scene understanding tasks like semantic segmentation. In this work, we focused on standalone models trained with conventional image-level augmentations, leaving this integration as a promising avenue for further research.

### v3

@@ 7 Discussion and Conclusion
This work presented an extensive analysis of the robustness of segmentation models under localized corruptions, considering both natural and adversarial corruptions. We first introduced metrics that extend the classic pixel-wise accuracy, commonly used in segmentation, to provide a more comprehensive analysis of spatial robustness, both within and beyond the corrupted regions. We then proposed a framework that enables a systematic evaluation of localized natural and adversarial corruptions by allowing the selection of parameters, including corruption ratios and model patch sizes.
Our analysis of adversarial perturbations also improves upon existing strategies in the literature by demonstrating that, in dense prediction tasks such as semantic segmentation, targeting multiple pixels introduces a complex multi-objective problem. This complexity makes it difficult to assess robustness in the worst-case adversarial scenario using a single localized perturbation. Therefore, we proposed the region-aware multi-attack adversarial analysis, a method to assess the spatial robustness of models by applying multiple $\infty$-norm attacks while continuously adjusting the fooling area.
All findings were evaluated across a diverse set of segmentation models in a driving scenario, specifically using the Cityscapes dataset [cordts2016cityscapes]. Our results revealed distinct differences in model behavior under natural and adversarial corruptions. Furthermore, we highlighted the importance of finding a trade-off between these two robustness aspects, which are not necessarily correlated. As a preliminary step toward addressing this challenge, we explored ensemble strategies that combine convolution-based models with transformer-based architectures.
In a future work, we aim at further investigating this trade-off by exploring more complex ensembling approaches, extending the analysis to other tasks and types or forms of corruptions, and providing theoretical insights into this intriguing mixed property. Another important direction is to integrate the proposed localized robustness analysis into model training strategies through the use of localized augmentation techniques (e.g., Cutout [devries2017improved_cutout] and CutMix [yun2019cutmix]), which are rarely applied in complex scene understanding tasks such as semantic segmentation. This could provide valuable insights into the development of novel defense mechanisms and robust training approaches. In this work, we focused on standalone models trained with conventional image-level augmentations, leaving this integration as a promising direction for future research.
@@
Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the MUR (Ministero dell’Università e della Ricerca) National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.